OpenStudy (anonymous):

six times the quantity , x-2, is 24. find the number , x.

OpenStudy (anonymous):

6(x-2)=24 6x-12=24 6x=24+12 6x=36 x=6

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

;)